The length of a rectangle is 3 cm more than twice its width. If the perimeter of the rectangle is 42 cm, find the length of the rectangle.

Answer by Cromlix(4381)

Hi there,
Make the width = 'x'
Width = x
Length = 3 + 2x
Perimeter = 2 x Width + 2 x Length
42 = 2(x) + 2(3 + 2x)
42 = 2x + 6 + 4x
Collect like terms
2x + 4x = 42 - 6
6x = 36
x = 6
Width = 6cm
Length = 15cm
Hope this helps:-)